# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v2

@@ Abstract
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 10 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
@@ I Introduction
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. EMT has found successful applications in various domains, including engineering scheduling [5, 8], nonlinear equation systems [9], feature selection in machine learning [10], anomaly detection [11], point cloud registration [12, 13], and reinforcement learning [14].
@@
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. These distinctive requirements pose unsolvable challenges for MTEA implementation in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45]. Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. The main contributions of MToP are summarized as follows:
@@
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types (see Table IV). Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types (see Table V). In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications (see Table VII). Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities (see Table VIII).
@@ II-A1 Test Module
Figure 1: Examples of graphic display in the test module of MToP.
The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms. It facilitates running the selected algorithm on the chosen problem within MToP. The module includes visualization functions for both single-objective and multi-objective MTO problems. In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig. 1 (a)-(b). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig. 1 (c). Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig. 1 (d). Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig. 1 (e).
@@
Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space with the number of generations.
(a) Gen=1
(b) Gen=100
(c) Gen=200
(d) Gen=500
(e) Gen=700
(f) Gen=1000
Figure 3: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the decision space with the number of generations.
In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior. Figures 2 and 3 illustrate an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space. Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=100,200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500,700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task. However, these two tasks do not precisely overlap in the decision space. This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
@@
The workflow of API calls for executing an algorithm on a problem is depicted in Fig. 4. The process initiates with the Start_Button event in MTO_GUI.m, triggering the parameter setting API Set_Parameter within the selected algorithm MFEA.m. Subsequently, it proceeds to the Run function of the algorithm, which initially invokes the Initialization API of the algorithm base class Algorithm.m, and then iterates through the function evaluation API Evaluation. During the initialization phase, the parameter settings of the problem CEC17_MTP1.m to be solved are obtained via the problem base class Problem.m. Similarly, the Evaluation function retrieves the specific optimization task object through the problem base class. Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function Obj.m via the Call_Metric API.
@@ II-A2 Experiment Module
Figure 5: schematic plotting by the experiment module of MToP.
@@ II-A3 Data Process Module
Each unique execution within the experiment module is considered a data object and is saved to a data file named MTOData.mat. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II. This saved data can subsequently be reloaded by the experiment module of MToP.
@@ II-B File Structure
Figure 6: File structure of MToP.
In the root directory MTO/ of MToP, you’ll find a script file named mto.m and four folders: Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is depicted in Fig. 6.
The Algorithms/ folder contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m. Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders. The Algorithms/Utils/ folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
The Problems/ folder houses various categories of problems, the base function folder, and the problem base class Problem.m. Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders. Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
The Metrics/ folder contains all result evaluation metrics, organized into subfolders for single-objective and multi-objective optimization (Metrics/Single-objective/ and Metrics/Multi-objective). Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
Lastly, the GUI/ folder houses all files used by the GUI of MToP. Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
@@ II-C1 Algorithm
All algorithms within MToP inherit from the algorithm base class Algorithm.m. The properties and methods of this base class are detailed in Table III. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table IV. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table V.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
In Listing LABEL:list:algo, the variable pops is of cell data type, housing multiple populations as popspopulation1, population2, population3…. Each individual within pops is an Individual object, containing properties such as objective value (Obj), constraint violation (CV), and decision variables (Dec). Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the population variable, directly containing individuals as population(individual1:factor1, individual2:factor1, individual3:factor2…), along with an additional property (MFFactor).
@@
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables (Dec) of offspring individuals and is tailored to specific algorithms. Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object (Prob) for offspring fitness evaluation. This entails calculating the objective value (Obj) and constraint violation (CV) for all individuals. Following each loop iteration, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
TABLE IV: Multitask evolutionary algorithms included in MToP (current version).
Algorithm Year Publication Objective Task Framework Constraint Description
MFEA [19] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
LDA-MFEA [46] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
MFDE&MFPSO [47] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
G-MFEA [48] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
MFEA-GHS [49] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
MFEA-DV [50] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
MFEA-II [20] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
SREMTO [51] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
TLTLA [52] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
MFEA-AKT [21] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
ASCMFDE [53] 2021 TEVC Single Multi Multifactorial Yes Intertask transfer in aligned subspace for MFDE
AT-MFEA [26] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
MFEA-DGD [54] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
IMEA [55] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
EMEA [24] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
MFMP [22] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
DEORA [33] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
MKTDE [56] 2022 TEVC Single Multi Multi-population Yes Meta-knowledge transfer-based multitask DE
MTES [57] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
BLKT-DE [58] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
MTSRA [44] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
MTEA-AD [23] 2022 TEVC Single Multi/Many Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
MTES-KG [59] 2023 TEVC Single Multi/Many Multi-population Yes Multitask ES with knowledge-guided external sampling
BoKT [30] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
TRADE [60] 2023 TCYB Single Many Multi-population Yes Transferable adaptive parameter DE for many-task optimization
MTEA-SaO [61] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
KR-MTEA [62] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive multi-task EA based on knowledge reuse
SBO [41] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
MaTDE [40] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
EMaTO-MKT [32] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
MO-MFEA [42] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominated sort and crowding distance
MO-MFEA-II [63] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
EMT-PD [64] 2021 TSMC-S Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
EMT-GS [29] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
MM-DE [65] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
AMT-NSGA-II [66] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
EMT-ET [34] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
MOMFEA-SADE [67] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
MTEA-D-DN [68] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
MTDE-MKTA [69] 2024 TEVC Multi Multi Multi-population Yes Multitask DE with multiple knowledge types and transfer adaptation
TABLE V: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
Algorithms Year Publication Objective Category Constraint Description
GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
jDE [70] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
JADE [71] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
rank-DE [72] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
SHADE [73] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
jSO [74] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
LSHADE44 [75] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
L-SHADE [76] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
CAL-SHADE [77] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
NL-SHADE-RSP [78] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
PSO [79] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
CSO [80] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
MPA [81] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
EO [82] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
AO [83] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
CMA-ES [84] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
IPOP-CMA-ES [85] 2005 CEC Single Evolution Strategy Yes Restart CMA-ES with increasing population size
sep-CMA-ES [86] 2008 PPSN Single Evolution Strategy Yes Separable CMA-ES for large-scale optimization
xNES-as [87] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
R1-NES [88] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
xNES&SNES [89] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
OpenAI-ES [90] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
DES [91] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
MTV-DE [92] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
ECHT [93] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
FROFI [94] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
C2oDE [95] 2019 TSMC-S Single Constraint Handling Yes Composite DE for constrained optimization
CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
DeCODE [96] 2021 TSMC-S Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
VMCH [97] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
CEDE-DRL [98] 2023 SWEC Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
CCEF-ECHT [99] 2024 TSMC-S Single Constraint Handling Yes Competitive and cooperative ensemble of constraint handling techniques
SPEA2 [100] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
NSGA-II [101] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
MO-CMA-ES [102] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
MOEA/D-DE [103] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
LMOCSO [104] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
CMOCSO [105] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization

### v3

@@ Abstract
Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years. It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques. Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 20 performance metrics. Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems. MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization. More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
@@ I Introduction
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task. For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space [8] as follows:
$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$. During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision. EMT has found successful applications in various domains, including engineering scheduling [5, 9], nonlinear equation systems [10], feature selection in machine learning [11], anomaly detection [12], point cloud registration [13, 14], and reinforcement learning [15].
@@
However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks. Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques. Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics. Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem. These bring uncertainty to the pattern of results display and analysis. These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms. Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45]. Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications. The reason for using MATLAB is that it is a widely used programming language in the field of EC. Moreover, MATLAB provides a user-friendly graphical interface and a powerful numerical computing environment, making it an ideal choice for developing MToP. More importantly, the existing MTEAs and MTO problems are mainly implemented in MATLAB, which can be directly integrated into MToP. The main contributions of MToP are summarized as follows:
@@
    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API. Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types. Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types. In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications. Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
@@ II Architecture of MToP
MToP consists of key modules including the test, experiment, and data processing modules, which facilitate algorithm execution, visualization, and data analysis. These components allow for flexible experimentation through both GUI and command line interfaces. Following standardized code patterns, MToP integrates various algorithms and performance metrics.
@@ II-A1 Test Module
Figure 1: Examples of graphical display in the test module of MToP. (a) and (b) illustrate the landscapes of single-objective problems with different tasks in the one- and two-dimensional unified search space, respectively. (c) shows the feasible regions of a single-objective problem with different tasks in the two-dimensional unified search space. (d) depicts the Pareto front of a multi-objective problem with multiple tasks. (e) displays the convergence behavior of metrics after executing algorithms on problems.
The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms. It facilitates running the selected algorithm on the chosen problem within MToP. The module includes visualization functions for both single-objective and multi-objective MTO problems. Users can select different options to display depending on the problem type, resulting in different visual outputs, as demonstrated below. In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig. 1 (a)-(b). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig. 1 (c). Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig. 1 (d). Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig. 1 (e).
In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior. Figure 2 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space. CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives. The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$. The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.4$. Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
@@
(g) Gen=1
(h) Gen=100
(i) Gen=200
(j) Gen=500
(k) Gen=700
(l) Gen=1000
Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations. The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=100 to 200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task. However, these two tasks do not precisely overlap in the decision space. This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
@@
The workflow of API calls for executing an algorithm on a problem is depicted in Fig. 3. The process initiates with the <Start_Button> event in <MTO_GUI.m>, triggering the parameter setting API SetParameter() within the selected algorithm <MFEA.m>. Subsequently, it proceeds to the Run() function of the algorithm, which initially invokes the Initialization() API of the algorithm base class <Algorithm.m>, and then iterates through the function evaluation API Evaluation(). During the initialization phase, the parameter settings of the problem <CEC17_MTP1.m> to be solved are obtained via the problem base class <Problem.m>. Similarly, the Evaluation() function retrieves the specific optimization task object through the problem base class. Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function <Obj.m> via the CallMetric() API.
@@ II-A2 Experiment Module
Figure 4: schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
@@ II-A3 Data Process Module
Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>. The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II. This saved data can subsequently be reloaded by the experiment module of MToP.
@@
In practical use, the specific operations for splitting and merging data are as follows:
  • Data Splitting: If an experiment involves running both the MFEA and EMEA algorithms on same problems, the data can be split by selecting it in the <Data Tree> and clicking the <Algorithms Split> button to generate separate datasets for each algorithm. Similarly, data can also be split by independent runs or problems using the respective options.
  • Data Merging: If the MFEA and EMEA algorithms were run on same problems in separate experiments, the data can be merged by selecting both datasets and clicking the <Algorithms Merge> button, provided the settings are identical except for the algorithm. This process can also be used to merge datasets by independent runs or problems.
@@ II-B File Structure
Figure 5: File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
In the root directory <MTO/> of MToP, you’ll find a script file named <mto.m> and four folders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>. The structure of these folders is depicted in Fig. 5.
The <Algorithms/> folder contains various categories of algorithms, utility functions, and the algorithm base class <Algorithm.m>. Specific algorithm files such as <MFEA.m>, <MO_MFEA.m>, and <GA.m> are stored within their respective classification folders. The <Algorithms/Utils/> folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
The <Problems/> folder houses various categories of problems, the base function folder, and the problem base class <Problem.m>. Specific problem files like <WCCI20_MTSO1.m>, <CMT3.m>, and <CEC19_MaTMO6.m> are stored under their corresponding classification folders. Real-world application problem files are located in the <Real-World Application/> folder, while the <Problems/Base/> folder contains public base functions for problems.
The <Metrics/> folder contains all result evaluation metrics, organized into subfolders for single-objective <Metrics/Single-objective/> and multi-objective <Metrics/Multi-objective> optimization. Specific metric files include <Obj.m>, <Obj_MTS.m>, and <IGD.m>, among others.
Lastly, the <GUI/> folder houses all files used by the GUI of MToP. Among them, <MTO_GUI.m> serves as the main file of the GUI, while <MTO_CMD.m> provides functionality for executing experiments via the command line.
@@ II-C1 Algorithm
All algorithms within MToP inherit from the algorithm base class <Algorithm.m>. The properties and methods of this base class are detailed in Table III. This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself. The MTEAs implemented in MToP are outlined in Table LABEL:supp:tab:algo-mt of the supplementary files. Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems. Specific algorithms belonging to this category are listed in Table LABEL:supp:tab:algo-st.
Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively. The algorithm labels provided in the second line serve as identifiers for classification within the GUI. The run() function within each class orchestrates algorithm execution, with <Algo> representing the object itself and <Prob> denoting the problem object to be solved. Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
In Listing LABEL:list:algo, the variable <pops> is of <cell> data type, housing multiple populations as <population1, population2, population3...>. Each individual within <pops> is an <Individual> object, containing properties such as objective value <Obj>, constraint violation <CV>, and decision variables <Dec>. Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the <population> variable, directly containing individuals as <individual1:factor1, individual2:factor1, individual3:factor2...>, along with an additional property <MFFactor>.
@@
In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks. Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function. The Generation() function operates on the decision variables <Dec> of offspring individuals and is tailored to specific algorithms. Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object <Prob> for offspring fitness evaluation. This entails calculating the objective value <Obj> and constraint violation <CV> for all individuals. Following each loop iteration, the environmental selection function Selection() is invoked to update the new population. While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function. Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes. The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.